If x units of a product cost C dollars to manufacture and earn revenue of R dollars, the value of x where the expressions for C and R are equal are called the break-even quantity, the number of units that produce zero profit. Find the break-even quantity when C=15x+54000 and R=18x.
Thank you!

Answer by Theo(13342)

C = 15x + 54000
R = 18x
break even is when R = C
this means that:
18x = 15x + 54000
subtract 15x from both sides of this equation to get:
3x = 54000
divide both sides of this equation by 3 to get:
x = 18000
the breakeven point is when x = 18000
R = 18x = 18 * 18000 = 324000
C = 15x + 54000 = 15 * 18000 + 54000 = 270000b + 54000 = 324000